OpenStudy (anonymous):

I have a question: I'm trying to evaluate a limit but am having a tough time using the limit laws: lim ( sqrt(121+h) - 11)/ h h->0

myininaya (myininaya):

multiplu top and bottom by the conjugate of the top

myininaya (myininaya):

\[\frac{\sqrt{121+h}-11}{h}*\frac{\sqrt{121+h}+11}{\sqrt{121+h}+11}\] \[\frac{121+h-121}{h(\sqrt{121+h}+1}\]

myininaya (myininaya):

\[\frac{1}{\sqrt{121+h}+11}\]

myininaya (myininaya):

now nut plug in 0

myininaya (myininaya):

lol just*

myininaya (myininaya):

nut plug lol

OpenStudy (anonymous):

hehehe

myininaya (myininaya):

i left my 1) off above do you see that?

myininaya (myininaya):

the last fraction i wrote is right but the one before that is missing 1)

OpenStudy (anonymous):

Okay

myininaya (myininaya):

so you get 1/22

OpenStudy (anonymous):

Oh wait ... duh, you're right ..